OpenStudy (anonymous):

Write a formula for the nth term of the given geometric sequence. Do not use a recursion formula. 2, ½, 1/8, 1/32

OpenStudy (anonymous):

nth term = a (r)^(n-1) where r is the common ratio n a is the first series for the given series a=2 and n=1/4 nth term=2(1/4)^(n-1)

OpenStudy (anonymous):

thats the answer?

OpenStudy (anonymous):

yes:)

OpenStudy (anonymous):

wow, that was fast ;)